Find the radius and center of each circle.

12. (x - 2)^2 + (y - 3)^2 = 16
13. x^2 + (y + 4)^2 = 8
14. (x + 1)^2 + (y - 2)^2 = 12
15. (x - 6)^2 + y^2 = 25

Answer by anantha(86)

sol:
the equation is of the form
(x-h)^2+(y-k)^2=r^2
then the center=(h,k)
and radius=r
1)
(x-2)^2+(y-3)^2=16
comparing with
(x-h)^2+(y-k)^2=r2
here h=2,k=3,r^2=16
center=(h,k)
center=(2,3)
r^2=16,r=sqrt(16)=4
radius=4
2)
(x+0)^2+(y+4)^2=8
here h=0,k=-4
center=(0,-4)
radius=sqrt(8)